WAVE-PACKET AND DENSITY MATRIX ANALYSIS OF COHERENT QUANTUM CONTROL; ”LASER MUSIC”

\underline{VADIM V. LOZOVOY}, IGOR PASTIRK, and MARCOS DANTUS, Department of Chemistry, Michigan State University, East Lansing, 48824.

A wave-packet and density matrix analysis is presented on some of the most successful techniques being used for laser control. We will present how three pulse four wave mixing can be used to combine the pump-dump method with quantum interference, multiphoton excitation and pulse shaping.